# Comparative Evaluation of 3D Reconstruction Methods for Object Pose Estimation

### v1

@@ 3 Benchmark
Some of the original YCB-V [111] objects were discontinued so we use the alternatives recommended by the YCB benchmark [5]. Thus, the objects collected in this benchmark may differ in texture and scale, with an average scale change of 4% of the object’s dimensions. Given the small object’s dimensions, such changes often fall below the distance thresholds used in the metrics. Therefore, we choose to keep the scale of the collected objects to be consistent with the current object’s reality.
Implementation Details.  The Nerfstudio [100] and SdfStudio [115] codebases are used to run Nerfacto [100], MonoSDF [116], VolSDF [113], UniSURF [75], Neus [109], and BakedSDF [114]. We use the author’s codebase for all other methods and for the pose estimation [55, 12], the MVS-texturing [106], the reconstruction evaluation [89] and the pose evaluation [3]. The mesh registration and post-processing are run with Open3D [118]. All methods run on a single NVIDIA A100 GPU except for RealityCapture that runs on a GeForce RTX 3060.
@@ 5 Experimental Evaluation
We recall the metrics to evaluate the reconstructions and the object poses estimated with those. The baseline is made of poses estimated with the original CAD models (Oracle). We then analyze how the quality of the estimated poses is impacted by the object properties, the texturing, and the amount of images used in the reconstruction.
Figure 1: 3D Reconstruction Evaluation. Left: Most reconstructions exhibit high geometric quality. Center: 3D reconstructions evaluated with the Average Recall (AR) of Megapose [55] (dashed and dotted) and FoundationPose [12] (uniform) when replacing the CAD models with the reconstructions. The pose evaluation is more discriminative than the geometric one (left) and the relative ranking of the methods is consistent between the two pose estimators. Right: Runtime, mesh appearance quality and pose quality [55]. Nerfacto [100] and Reality Capture [82] achieve the best peformances.
@@
The Visible Surface Discrepancy (VSD) measures the distance between the surfaces of the object at the ground-truth and estimated poses. The Maximum Symmetry-Aware Surface Distance (MSSD) measures the maximum surface distance between the two poses. This metric is symmetry-aware, i.e., it does not penalize the estimated pose for ambiguous symmetries. The Maximum Symmetry-Aware Projection Distance (MSPD) measures the visual discrepancy in pixel distance induced by an incorrect pose when rendering the object. The fraction of images for which the errors are below a given threshold defines the recall. It is averaged over several error thresholds and over images to form the AR for each error. The resulting three ARs are averaged to form the global AR that we report next. Interpretation: the VSD and MSSD are good indicators of the object pose quality for navigation or object-manipulation applications while the MSPD is better suited for applications that exploit object rendering, such as Augmented and Virtual Reality.
@@
Reconstruction and Pose Scores.  Fig. 1 shows the geometric quality of the reconstructions (left), their performance when integrated into FoundationPose [12] and Megapose [55] (center), and the correlation between the estimated poses [55], the appearance quality of the mesh, and the reconstruction runtime (right). The results demonstrate the practical advantage of the MVS-based Reality Capture [82] that achieves SotA reconstruction performance under the shortest runtime (recall that RealityCapture is run on a weaker GPU). Overall, there remains a 10-20% gap in quality between CAD models and the best 3D reconstructions (Fig. 1-center). Given that the relative pose performances of the reconstructions are consistent between the two pose estimators, one can expect advances in 3D reconstruction to close the gap for other pose estimators.
The reconstructions best for pose estimation usually exhibit high geometric and appearance quality. Still, there is no direct correlation between the reconstruction metrics and the pose scores: for example the geometric completeness of UniSURF [75] and iNGP [71] is relatively low, which is reflected in the pose performance. At the same time, RealityCapture [82] performs comparably to BakedSDF [114] in terms of completeness and accuracy but leads to slightly worse poses. Conversely, the reconstructed geometry of Nerfacto [100] is worse than MonoSDF [116] but their pose evaluation is on par. One interesting observation is that the geometric evaluation (Fig. 1-left) does not discriminate between VolSDF and MonoSDF (superimposed red and cyan curves) whereas the pose evaluation does. This validates the proposed reconstruction evaluation that not only measures the accuracy of the reconstruction itself but also the quality of the pose estimation it enables.
There is also no immediate correlation between the quality of the mesh’s appearance and the pose quality (Fig. 1-right). For example, Reality Capture reaches a higher PSNR than BakedSDF yet the latter achieves better pose performances. As for Colmap [87] and NeuralAngelo [59], they exhibit similar PSNR but the first leads to much worse pose estimation.
@@
Finer Pose Evaluation.  The BOP benchmark defines multiple metrics to assess how suitable the pose estimator, hence the 3D reconstruction, is for a given application such as augmented and virtual reality (MSPD), navigation (VSD), and object manipulation (MSSD). Due to the page limit, we report detailed results in Supp-C1. The VSD recall remains close to the baseline so objects are overall well-positioned, which is useful for global object reasoning such as object avoidance and navigation. However, there is a drop in MSSD recall that is particularly sensitive to small misplacements when objects exhibit high curvature regions. This is typical of grasping points so this signals a limitation when using reconstructed meshes for object manipulation. Another limitation relates to virtual reality because of the MSPD recall drop that indicates visual dissonance when rendering objects.
Object Properties. We split the YCB-V [111] objects into 8 groups based on their shape, texture, material properties, and the degree of change between the collected objects and the original ones. We show the results for 3 groups in Fig. 2 and the rest in Supp-C.3. The pose performance variations across groups are telling of some of the current 3D reconstruction challenges.
@@
Conversely, the reconstruction of objects with shiny or small texture elements performs worse than other categories, independently of the pose estimators. Several factors contribute to this drop. First, several of the objects in these categories have been discontinued or got their texture updated. The texture inconsistency between the test images and our reconstructions impedes the pose estimation, which is confirmed by the higher pose performance on the ‘legacy’ objects‘ than on the updated ones. Another factor is that small and shiny textures are harder to generate than Lambertian ones. When it comes to small objects ($\leq$ 8cm), the reconstruction generates a low-resolution texture, which decreases the amount of visual information relevant to disambiguating the estimated poses. This is all the more critical as those objects exhibit symmetries that can only be differentiated through their texture (07-pudding-box, 06-tuna-fish-can). As for shiny objects, the light reflection induces a visual discrepancy between the target texture and the reconstructed one.
@@
Number of images. Besides the quality of the pose estimation, an important property of the reconstruction methods is their data requirement. Fig. 8-left shows the geometric and Megapose evaluation [55] of reconstructions generated from varying numbers of images sampled uniformly around the object [33]. As expected, the results get better with the number of images and 75-100 images is enough to generate reconstructions close to the best ones. However, we observe that using sparse views can hinder the convergence of some reconstructions. We also observe that the performance does not increase linearly with the data which we attribute to non-uniform noise in the camera poses. These results show that another avenue for improvement includes robustness to noisy calibration and reconstructions from limited views. In parallel, qualitative results in Fig. 8-right illustrate examples where reconstructions are visually subpar yet informative enough for satisfying pose estimation. This suggests that incomplete meshes can be relevant for object pose estimation as long as the discriminative parts of the objects are reconstructed such as the outer edge of the bowl.
@@ 6 Conclusions
In this work, we have considered the problem of reconstructing 3D models from images such that they can be used for object pose estimation. In particular, we aimed to answer the question of whether state-of-the-art 3D reconstruction algorithms can be readily used to remove the need for CAD models during training and testing of object pose estimators. We have created a new benchmark, based on the YCB-V [17, 111] dataset, that can be used to measure the impact of replacing CAD models with reconstructed meshes. We have evaluated multiple state-of-the-art reconstruction methods and our results show that reconstructed models or their implicit render can often readily replace CAD models. Still, there remains a considerable gap between CAD models and image-based meshes for certain types of objects that are hard to reconstruct, e.g., small and / or shiny objects.
Several interesting observations can be drawn from our results: (1) The classical, handcrafted Reality Capture often performs similarly to state-of-the-art learning-based approaches. It is also very efficient and able to reconstruct scenes within a few minutes, which makes it a strong baseline for future work. (2) The ranking between different 3D reconstruction approaches is quite consistent between the two evaluated pose estimators. There thus is hope that all pose estimators will automatically benefit from advances in the area of 3D reconstruction. (3) Incomplete reconstructions, where parts of the object cannot be reconstructed, are not necessarily problematic as the resulting models can still enable accurate pose estimation as long as enough parts are reconstructed. This observation validates our idea of evaluating reconstruction systems inside a higher-level task as classical reconstruction-based metrics, e.g., accuracy and completeness, are falsely indicating the resulting models as low-quality reconstructions. (4) There is not a clear relation between the quality of the estimated 3D scene geometry and pose accuracy. For example, the renderings produced directly by the iNGP network achieve quite good performance with Megapose whereas the renderings of the iNGP mesh fall behind. This is a promising result as it suggests that objects hard to reconstruct can still allow accurate pose estimation if the novel view synthesis performs well. (5) In practice, one would ideally use as few images as possible for the reconstruction to minimize both capture and reconstruction time. Our results indicate room for improvement in the few-image regime.
@@
Acknowledgements
This work was supported by the Czech Science Foundation (GACR) EXPRO (grant no. 23-07973X). This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254). The authors would like to acknowledge the “Research and Innovation Centre on Advanced Industrial Production (RICAIP)“ supported by European Structural and Investment Funds and Operational Programme Research, Development and Education via the Ministry of Education, Youth and Sports of the Czech Republic (ID: CZ.02.1.01/0.0/0.0/17_043/0010085) and from the European Union’s Horizon 2020 research and innovation program under grant agreement 857306. This work was supported by the Grant Agency of the Czech Technical University in Prague, grant No.SGS23/172/OHK3/3T/13. This work was co-funded by European Union under the project Robotics and advanced industrial production - ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590).
Supplementary: Comparative Evaluation of 3D Reconstruction Methods for Object Pose Estimation
The supplementary material is organized as follows: Sec. A describes the data acquisition and the information related to the release as announced in Sec.3 of the main paper. Sec. B provides implementation details relative to the reconstructions, the pose estimation, and the evaluation, as announced in Sec. 4 of the main paper. Sec.C reports results complementary to Sec. 5 of the main paper.
Appendix A Dataset
The proposed benchmark answers the question to what degree 3D models reconstructed by current 3D reconstruction algorithms can replace the CAD models commonly used for object pose estimation. To this end, we compare the performance of two SotA object pose estimators, FoundationPose and Megapose [55, 12], when using the 3D reconstructed models against the use of classical CAD models. Thus, we mainly measure 3D reconstruction performance by the accuracy of the estimated poses rather than the accuracy of the resulting 3D models itself.
@@ Appendix A Dataset
In this section, we describe how the images in (1) were collected (Sec. A.1) and how the data is released (Sec. A.2).
A.1 Data Collection Setup
@@
The camera is calibrated using open-source solutions: the OpenCV routine for the camera calibration [15] and the MoveIt [20] routine for the hand-eye calibration  [^2] https://github.com/ros-planning/moveit_calibration. Afterward, the camera can be positioned with an error of up to 5mm in translation and 0.1 in rotation. The resulting camera calibration is used to undistort the images using the open-source COLMAP [87] library.
@@
Pose Refinement.  The camera extrinsics of the collected images are obtained from the kinematic chain of the robot that can be relatively noisy. One source of such noise is in the robotic arm’s positional encoders: they are known to be noisier in configurations that are close to degenerate, e.g., when the arm is extended. This can affect the quality of the reconstruction methods so we refine the camera extrinsics in two steps: i) we use sequentially the pose refinement capabilities of iNGP [71] that update the poses consistently with the trained 3D reconstruction; ii) we then use the iNGP-refined poses to generate a sparse 3D reconstruction and refine the camera poses again with bundle adjustment using COLMAP [87]. After this step, the camera extrinsics remain fixed.
Image Registration.  The YCB-V dataset [111] provides test images depicting the objects and their ground-truth poses. These poses are the rotation and the translation between the camera coordinate frame and the object’s coordinate frame, which is the coordinate frame of the original CAD model. This coordinate frame is different from the ‘world’ coordinate frame where we collect the images and reconstruct the objects. Thus, the YCB-V ground-truth poses are not compatible with the coordinate frame of the 3D reconstructions. Consequently, we register the collected images to the original CAD model for each object. To do so, we first generate 3D reconstructions for all objects in their coordinate frame, and then estimate the rigid transform between the reconstruction’s coordinate frame and the CAD model’s coordinate frame with ICP [10, 118]. The ICP is initialized with user-defined coarse alignments.
@@
A.2 Data Release
The released data is available at
The webpage contains instructions on how to download the data from the Apache server, either through a web page or a simple command line tool, e.g., wget.
Released Data and Format.  We release the calibrated undistorted images of the 21 YCB-V [111] objects collected in Sec. A.1, the objects’ masks, and the meshes generated by the reconstruction methods evaluated in the benchmark.
For each object, the undistorted images, their extrinsics, the camera intrinsics, and the object masks are packed in a single zip file. The images and the masks are in png format, the extrinsics and intrinsics are released under both the Nerfstudio [100] camera pose convention 333docs.nerf.studio/quickstart/data_conventions.html and the COLMAP [87] camera pose convention 444colmap.github.io/format.html#text-format These are two of the most commonly used camera pose conventions for 3D reconstruction. Dataparsers for these two formats are available at the respective repositories.

### v2

@@ 3 Dataset
Some of the original YCB-V [112] objects were discontinued so we use the look-alike alternatives recommended by the YCB benchmark [5]. Thus, the objects collected in this benchmark may differ in texture and scale, with an average scale change of 4% of the object’s dimensions. Given the small object’s dimensions, such changes often fall below the distance thresholds used in the metrics. Therefore, we choose to keep the scale of the collected objects to be consistent with the current object’s reality.
@@ 5 Experimental Evaluation
We compare the performance of SotA object pose estimators [55, 110] when using the 3D reconstructed models against the use of classical CAD models. We follow the object pose evaluation guidelines defined in the BOP benchmark [45, 46] that provides accurate CAD models. These CAD models are used as a reference for the reconstruction evaluation and to estimate the baseline object poses (Oracle). We recall the metrics used to evaluate the reconstructions and the object poses. We then analyze how the quality of the estimated poses is impacted by the 3D reconstructions, the object properties, the texturing, and the amount of images used in the reconstruction.
Figure 1: 3D Reconstruction Evaluation. Left: Most reconstructions exhibit high geometric quality. Center: 3D reconstructions evaluated with the Average Recall (AR) of Megapose [55] (dashed and dotted) and FoundationPose [110] (uniform) when replacing the CAD models with the reconstructions. The pose evaluation is more discriminative than the geometric one (left) and the relative ranking of the methods is consistent between the two pose estimators. Right: Runtime, mesh appearance quality and pose quality [55]. Nerfacto [100] and Reality Capture [82] achieve the best performances when accounting for both the pose recall and the runtimes.
Implementation Details.  The Nerfstudio [100] and SdfStudio [116] codebases are used to run Nerfacto [100], MonoSDF [117], VolSDF [113], UniSURF [75], NeuS [109], and BakedSDF [114]. We use the author’s codebase for all other methods and for the pose estimation [55, 110], the MVS-texturing [106], the reconstruction evaluation [90] and the pose evaluation [3]. The mesh registration and post-processing are run with Open3D [119]. All methods run on a single NVIDIA A100 GPU except for RealityCapture that runs on a GeForce RTX 3060 for it requires a Windows installation.
@@
The Visible Surface Discrepancy (VSD) measures the distance between the surfaces of the object at the ground-truth and estimated poses. The Maximum Symmetry-Aware Surface Distance (MSSD) measures the maximum surface distance between the two poses. This metric is symmetry-aware, i.e., it does not penalize the estimated pose for ambiguous symmetries. The Maximum Symmetry-Aware Projection Distance (MSPD) measures the visual discrepancy in pixel distance induced by an incorrect pose when rendering the object. The fraction of images for which the errors are below a given threshold defines the recall. It is averaged over several error thresholds and over images to form the AR for each error. The resulting three ARs are averaged to form the global AR that we report next. Interpretation: the VSD and MSSD are good indicators of the object pose quality for navigation or object-manipulation applications while the MSPD is better suited for applications that exploit object rendering, such as Augmented and Virtual Reality [45].
Reconstruction and Pose Scores.  Fig. 1 shows quantitative results on the geometric quality of the reconstructions (left), their performance when integrated into FoundationPose [110] and Megapose [55] (center), and the correlation between the estimated poses [55], the appearance quality of the mesh, and the reconstruction runtime (right). The results demonstrate the practical advantage of the MVS-based Reality Capture [82] that achieves SotA reconstruction performance under the shortest runtime (recall that RealityCapture is run on a weaker GPU). Overall, there remains a 10-20% gap in quality between CAD models and the best 3D reconstructions (Fig. 1-center). Given that the relative pose performances of the reconstructions are consistent between the two pose estimators, one can expect advances in 3D reconstruction to close the gap for other pose estimators.
@@
The reconstructions best for pose estimation usually exhibit high geometric and appearance quality. Still, there is no direct correlation between the reconstruction metrics and the pose scores: for example the geometric completeness of UniSURF [75] and iNGP [71] is relatively low, which is reflected in the pose performance. At the same time, RealityCapture [82] performs comparably to BakedSDF [114] in terms of completeness and accuracy but leads to slightly worse poses. Conversely, the reconstructed geometry of Nerfacto [100] is worse than MonoSDF [117] but their pose evaluation is on par. One interesting observation is that the geometric evaluation (Fig. 1-left) does not discriminate between VolSDF and MonoSDF (superimposed red and cyan curves) whereas the pose evaluation does. This supports the necessity for the proposed benchmark and validates the proposed reconstruction evaluation that not only measures the accuracy of the reconstruction itself but also the quality of the pose estimation it enables.
There is also no immediate correlation between the quality of the mesh’s appearance and the pose quality (Fig. 1-right). For example, Reality Capture reaches a higher PSNR than BakedSDF yet the latter leads to better poses. As for Colmap [87] and NeuralAngelo [58], they exhibit similar PSNR but the former leads to much worse pose estimation.
@@
Finer Pose Evaluation.  The BOP benchmark defines multiple metrics to assess how suitable the pose estimator, hence the 3D reconstruction, is for a given application such as augmented and virtual reality (MSPD), navigation (VSD), and object manipulation (MSSD). Due to the page limit, we report the plots in Sec. C.1 of the supp. mat. The VSD recall of the reconstructions remains close to the baseline so objects are overall well-positioned, which is useful for global object reasoning such as object avoidance and navigation. However, there is a drop in MSSD recall that is particularly sensitive to small misplacements when objects exhibit high curvature regions. This is typical of grasping points so this signals a limitation when using reconstructed meshes for object manipulation. Another limitation relates to virtual reality because of the MSPD recall drop that indicates visual dissonance when rendering objects.
Object Properties. We split the YCB-V [112] objects into 8 groups based on their shape, texture, material properties, and the degree of change between the collected objects and the original ones. We show the plots for 3 groups in Fig. 2 and the rest in Sec. C.3 of the supp. mat. The pose performance variations across groups are telling of some of the current 3D reconstruction challenges.
@@
Conversely, the reconstruction of objects with shiny or small texture elements performs worse than other categories, independently of the pose estimators. Several factors contribute to this drop. First, several of the objects in these categories have been discontinued or got their texture updated. The texture inconsistency between the test images and our reconstructions impedes the pose estimation, which is confirmed by the higher pose performance on the ‘legacy’ objects, i.e., the objects that did not change, than on the updated ones. Another factor is that small and shiny textures are harder to generate than Lambertian ones. When it comes to small objects ($\leq$ 8cm), the reconstruction generates a low-resolution texture, which decreases the amount of visual information relevant to disambiguating the estimated poses. This is all the more critical as those objects exhibit symmetries that can only be differentiated through their texture (07-pudding-box, 06-tuna-fish-can). As for shiny objects, the light reflection induces a visual discrepancy between the target texture and the reconstructed one.
@@
Number of Images. Besides the quality of the pose estimation, an important property of the reconstruction methods is their data requirement. Fig. 8-left shows the geometric and Megapose evaluation [55] of reconstructions generated from varying numbers of images sampled uniformly around the object [32]. As expected, the results get better with the number of images and 75-100 images is enough to generate reconstructions close to the best ones. However, we observe that using sparse views can hinder the convergence of some reconstructions. We also observe that the performance does not increase linearly with the data which we attribute to non-uniform noise in the camera poses. One source of such noise is in the robotic arm’s positional encoders: they are known to be noisier in configurations that are close to degenerate, e.g., when the arm is extended. These results show that another avenue for improvement includes robustness to noisy poses and reconstructions from limited views. In parallel, qualitative results in Fig. 8-right illustrate examples where reconstructions are visually subpar yet informative enough for satisfying pose estimation: the two instances of the object lead to similar pose estimation performance. This suggests that incomplete meshes can be relevant for object pose estimation as long as the discriminative parts of the objects are reconstructed such as the outer edge of the bowl.
@@ 6 Conclusions
In this work, we have considered the problem of reconstructing 3D models from images such that they can be used for object pose estimation. In particular, we aimed to answer the question of whether state-of-the-art 3D reconstruction algorithms can be readily used to remove the need for CAD models during training and testing of object pose estimators. We have created a new benchmark, based on the YCB-V [15, 112] dataset, that can be used to measure the impact of replacing CAD models with reconstructed meshes. We have evaluated multiple state-of-the-art reconstruction methods. Our results show that reconstructed models or their renderings can often readily replace CAD models. Still, there is a considerable gap between CAD and image-based models for certain types of objects that are hard to reconstruct, e.g., small or shiny objects.
Several interesting observations can be drawn from our results: (1) The classical, handcrafted Reality Capture often performs similarly to state-of-the-art learning-based approaches. It is also very efficient and able to reconstruct scenes within a few minutes, which makes it a strong baseline for future work. (2) The ranking between different 3D reconstruction approaches is quite consistent between the two evaluated pose estimators. There thus is hope that all pose estimators will automatically benefit from advances in the area of 3D reconstruction. (3) Incomplete reconstructions, where parts of the object cannot be reconstructed, are not necessarily problematic as the resulting models can still enable accurate pose estimation as long as enough parts are reconstructed. This observation validates our idea of evaluating reconstruction systems inside a higher-level task as classical reconstruction-based metrics, e.g., accuracy and completeness, are falsely indicating the resulting models as low-quality reconstructions. (4) There is not a clear relation between the quality of the estimated 3D scene geometry and pose accuracy. For example, the renderings produced directly by the iNGP network achieve quite good performance with Megapose whereas the renderings of the iNGP mesh fall behind. This is a promising result as it suggests that objects that are hard to reconstruct can still allow accurate pose estimation if the novel view synthesis performs well. This also supports the necessity for the proposed benchmark that evaluates 3D reconstruction on their integration with pose estimators rather than with the standard reconstruction evaluations. (5) In practice, one would ideally use as few images as possible for the reconstruction to minimize both capture and reconstruction time. Our results indicate room for improvement in the few-image regime.
@@
Acknowledgements.  This work was supported by the Czech Science Foundation (GACR) EXPRO (grant no. 23-07973X), the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254), the Grant Agency of the CTU in Prague, grant No.SGS23/172/OHK3/3T/13, the EU under the project Robotics and advanced industrial production - ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590), the EU Horizon 2020 project RICAIP (grant agreement No 857306).
The supplementary material is organized as follows: Sec. 7 describes the data acquisition and the information related to the release as announced in Sec.3 of the main paper. Sec. 8 provides implementation details relative to the reconstructions, the pose estimation, and the evaluation, as announced in Sec. 4 of the main paper. Sec.9 reports further quantitative results that support the conclusions drawn in Sec. 5 of the main paper.
7 Dataset
The proposed benchmark answers the question “To what degree 3D models reconstructed by current 3D reconstruction algorithms can replace the CAD models commonly used for object pose estimation?” To this end, we compare the performance of two SotA object pose estimators, FoundationPose and Megapose [55, 110], when using the 3D reconstructed models against the use of classical CAD models. Thus, we mainly measure 3D reconstruction performance by the accuracy of the estimated poses rather than the accuracy of the resulting 3D models itself.
@@ 7 Dataset
In this section, we describe how the images in (1) were collected (Sec. 7.1) and how the data is released (Sec. 7.2).
7.1 Data Collection Setup
@@
The camera is calibrated using open-source solutions: the OpenCV routine for the camera calibration [13] and the MoveIt [19] routine for the hand-eye calibration [^3] https://github.com/ros-planning/moveit_calibration. Afterward, the camera can be positioned with an error of up to 5mm in translation and 0.1 in rotation. The resulting camera calibration is used to undistort the images using the open-source COLMAP [87] library. The camera poses obtained from the kinematic chain of the robot are later refined with an offline optimization (see the paragraph ‘Pose refinement’ below).
@@
Image Registration.  The YCB-V dataset [112] provides test images depicting the objects and their ground-truth poses. These poses are the rotation and the translation between the camera coordinate frame and the object’s coordinate frame, which is the coordinate frame of the original CAD model. This coordinate frame is different from the ‘world’ coordinate frame where we collect the images and reconstruct the objects. Thus, the YCB-V ground-truth poses are not compatible with the coordinate frame of the 3D reconstructions. Consequently, we register the collected images to the original CAD model for each object. To do so, we first generate 3D reconstructions for all objects in their coordinate frame with COLMAP [87] then estimate the rigid transform between the reconstruction’s coordinate frame and the CAD model’s coordinate frame with ICP [9, 119]. The ICP is initialized with user-defined coarse alignments.
Pose Refinement.  The camera extrinsics of the collected images are obtained from the kinematic chain of the robot that can be relatively noisy. One source of such noise is in the robotic arm’s positional encoders: they are known to be noisier in configurations that are close to degenerate, e.g., when the arm is extended. This can affect the quality of the reconstruction methods so we refine the camera extrinsics in two steps: i) we use sequentially the pose refinement capabilities of iNGP [71] that update the poses consistently with the trained 3D reconstruction; ii) we then use the iNGP-refined poses to generate a sparse 3D reconstruction and refine the camera poses again with bundle adjustment using COLMAP [87]. After this step, the camera extrinsics remain fixed. Although some evaluated methods have pose refinement capabilities, we turn them off so that all reconstruction methods share the same coordinate frame. Without this, the reconstruction’s coordinate frames may be slightly different for each method depending on the amplitude of the pose refinement, which would introduce inconsistencies in the evaluation.
@@
7.2 Data Release
The released data is available on the project’s website:
The webpage contains instructions on how to download the data from an Apache server, either through a web page or a simple command line tool, e.g., wget.
Released Data and Format.  We release the calibrated undistorted images of the 21 YCB-V [112] objects collected in Sec. 7.1, the objects’ masks, and the meshes generated by the reconstruction methods evaluated in the benchmark.
For each object, the undistorted images, their extrinsics, the camera intrinsics, and the object masks are packed in a single zip file. The images and the masks are in png format, the extrinsics and intrinsics are released under both the Nerfstudio [100] camera pose convention444docs.nerf.studio/quickstart/data_conventions.html and the COLMAP [87] camera pose convention555colmap.github.io/format.html#text-format. These are two of the most commonly used camera pose conventions for 3D reconstruction. Dataparsers for these two formats are available at the respective repositories.